OpenStudy (anonymous):

What is the angle in radians between the vectors -7 + 8 + 4 and 1 + 2 + 3

OpenStudy (anonymous):

are these vectors V=-7i+8j+4k and U=i+2j+3k ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you take the dot product between them to get 21 and then you know that 21 is the lengths multiplied together and multiplied by the sin of the angle so sin(theta)=21/((7^2+8^2+4^2)(1+2^2+3^2))^1/2

OpenStudy (anonymous):

sorry its cos not sin

OpenStudy (anonymous):

ok thank you